OpenStudy (anonymous):

A rectangular prism is truncated removing 30% of its height. The original height of the prism was 100 cm and it has a length of 6 cm and a width of 12 cm. What is the volume of the truncated prism?

OpenStudy (anonymous):

Height is cut by 30% So new height became: \[New \; Height = H - \frac{30}{100}H = \frac{7}{10}H\]

OpenStudy (anonymous):

So: \[Volume = 6 \times 12 \times \frac{7}{10}H\] Previous volume was : \[100 = L \times B \times H\] \[H = \frac{100}{72}\]

OpenStudy (anonymous):

Put it in the new volume we got..

OpenStudy (anonymous):

\[Volume = 6 \times 12 \times \frac{7}{10}H \implies 72 \times \frac{7 \times 100}{ 10 \times 72} \implies 70\]

OpenStudy (anonymous):

This stuff is so rough.

OpenStudy (anonymous):

Getting @Mastercat

OpenStudy (anonymous):

Rough ?? But I have used equation editor instead of drawing.. Ha ha ha..

OpenStudy (anonymous):

Let us start from the start...

OpenStudy (anonymous):

I miss algebra... so much easier.

OpenStudy (anonymous):

Volume given was : 100 L = 6 and H = 12 Use volume formula and find previous H.. \[Volume = 6 \times 12 \times H \implies 100 = 72 \times H\] Find H from it by dividing 72 both the sides.. \[H = \frac{100}{72}\] Getting till here @Mastercat

OpenStudy (anonymous):

Sorry W = 12 and not H in the first line I have written wrong...

OpenStudy (anonymous):

Now this we got was previous Height that is 100/72.. Now after truncating it gets reduced by 30%

OpenStudy (anonymous):

\[\frac{100}{72} - \frac{30}{100} \times \frac{100}{72} = \frac{100}{72} \times \frac{70}{100} \implies \frac{70}{72}\] This is the new height after truncation..

OpenStudy (anonymous):

Now again use the volume formula; \[Volume = 6 \times 12 \times \frac{70}{72} \implies 70\]

OpenStudy (anonymous):

But we have to find the volume of truncated portion and not the volume of prism that is left after truncation.. So, Subtract both the volumes: \[\large Volume(truncated) = 100 - 70 \implies \color{blue}{30}\]

OpenStudy (anonymous):

http://www.youtube.com/watch?v=bXScgXleLX8&feature=related = My Brain

OpenStudy (anonymous):

My internet always goes like an ant... I will see it some other day...

OpenStudy (anonymous):

Sorry... Just so complex. Im looking at it.